# Online Multivariate Regularized Distributional Regression for High-dimensional Probabilistic Electricity Price Forecasting

### v1

@@ 3 Forecasting Study
We propose modeling the multivariate distribution of the day-ahead electricity prices in increasing complexity. All models are updated online, i.e. using only the new data for each day. We differentiate between an adaptive estimation, which is updating a single, unconditional distributional parameter and the full conditional estimation linking the distribution parameter to explanatory variables.
@@
Model Mean / Location Marginal Distribution Dependence Structure
LARX Online conditional Adaptive but unconditional Adaptive but unconditional
Distr. Regression + Copula Online conditional Online conditional Adaptive but unconditional
Multivariate Dist. Regression Online conditional Online conditional Online conditional
Table 3: Increasingly complex model structures.
@@
Table 3 shows the increasing model complexity. We start with the established LARX models and naively estimate the unconditional residual distribution (denoted as $\operatorname{LARX}+\mathcal{N}(0,\sigma)$ and $\operatorname{LARX}+\mathcal{N}(0,\Sigma)$). We increase complexity by moving to full distributional model for the marginals and adding an adaptive estimation of the dependence structure using the Gaussian copula (denoted as oDistReg+Copula). Lastly, we estimate the full multivariate distribution in a conditional way using the proposed multivariate online distributional regression approach (denoted as $\mathcal{F}$, parameterization, method)). We describe the full model in the following and note that we additionally describe the hyper parameters in Appendix A.2. For all three complexity levels, we include a reference model that assumes independence to showcase the value-add of including the dependence structure. We model the mean/location for all regression models by
$g_{\mu}(\mu_{t,h})$ $=\beta_{\mu,0,h}+\sum_{l=1}^{L=7}+\beta_{\mu,l,h}y_{t-l,h}+\sum_{h\in\{0,...,23\}/h}\beta_{\mu,8+h,h}y_{t-1,h}$ (18)
$+\beta_{\mu,31,h}\operatorname{min}(\bm{\mathrm{y}}_{t-1})+\beta_{\mu,32,h}\operatorname{max}(\bm{\mathrm{y}}_{t-1})+\beta_{\mu,33,h}\operatorname{Q}_{10}(\bm{\mathrm{y}}_{t-1})+\beta_{\mu,34,h}\operatorname{Q}_{90}(\bm{\mathrm{y}}_{t-1})$
$+\beta_{\mu,35,h}\operatorname{Load}_{t,h}+\beta_{\mu,36,h}\operatorname{RES}_{t,h}+\beta_{\mu,37,h}\overline{\operatorname{Load}}_{t}+\beta_{\mu,38,h}\overline{\operatorname{RES}}_{t}$
$+\beta_{\mu,39,h}\operatorname{EUA}_{t}+\beta_{\mu,40,h}\operatorname{Gas}_{t}+\beta_{\mu,41,h}\operatorname{Coal}_{t}+\beta_{\mu,42,h}\operatorname{Oil}_{t}+\sum_{w}^{W=6}\beta_{\mu,42+w,h}\operatorname{WD}_{t,h}$
@@
$g_{\theta}(\theta_{t,h,h})$ $=\beta_{\theta,0,h}+\beta_{\theta,1,h}\operatorname{mean}(\bm{\mathrm{y}}_{t-1})+\beta_{\theta,2,h}\operatorname{SignedSquare}\left(\bm{\mathrm{\Sigma}}_{h,h}^{[t-1:t-7]}\right)+\beta_{\theta,3,h}\overline{\operatorname{Load}}_{t}+\beta_{\theta,4,h}\overline{\operatorname{RES}}_{t}$ (19)
$+\beta_{\theta,5,h}\operatorname{Load}_{t,h}+\beta_{\theta,6,h}\operatorname{RES}_{t,h}+\beta_{\theta,7,h}\operatorname{EUA}_{t}+\beta_{\theta,8,h}\operatorname{Gas}_{t}+\beta_{\theta,9,h}\operatorname{Coal}_{t}+\beta_{\theta,10,h}\operatorname{Oil}_{t}$
where $\operatorname{SignedSquare}(a)=\operatorname{sign}(a)\sqrt{|a|}$ is the signed square root and $\bm{\mathrm{\Sigma}}^{[t-1:t-7]}$ is the rolling empirical covariance matrix of $\bm{\mathrm{y}}_{t}$ for the last 7 days. For the LRA-based parameterization, we choose $r=2$ and model the elements of $\bm{\mathrm{V}}$ as
$g_{v}(v_{t,h,0})$ $=\beta_{v,0,h}+\beta_{v,1,h}\operatorname{mean}(\bm{\mathrm{y}}_{t-1})+\beta_{v,2,h}\operatorname{SignedSquare}\left(\bm{\mathrm{\Sigma}}_{h,h}^{[t-1:t-7]}\right)+\beta_{v,3,h}\operatorname{min}(\bm{\mathrm{y}}_{t-1})$ (20)
$+\beta_{v,4,h}\operatorname{max}(\bm{\mathrm{y}}_{t-1})+\beta_{v,5,h}\operatorname{Q}_{10}(\bm{\mathrm{y}}_{t-1})+\beta_{v,6,h}\operatorname{Q}_{90}(\bm{\mathrm{y}}_{t-1})+\beta_{v,7,h}\overline{\operatorname{Load}}_{t}+\beta_{v,8,h}\overline{\operatorname{RES}}_{t}$
$+\beta_{v,9,h}\operatorname{Load}_{t,h}+\beta_{v,10,h}\operatorname{RES}_{t,h}+\beta_{v,11,h}\operatorname{EUA}_{t}+\beta_{v,12,h}\operatorname{Gas}_{t}+\beta_{v,13,h}\operatorname{Coal}_{t}+\beta_{v,14,h}\operatorname{Oil}_{t}$
@@
$g_{\nu}(\nu_{t})$ $=\beta_{\nu,0}+\beta_{\nu,1}\operatorname{mean}(\bm{\mathrm{y}}_{t-1})+\sum_{w}^{W=6}\beta_{\nu,1+w,h}\operatorname{WD}_{t,h}+\beta_{\nu,8}\overline{\operatorname{Load}}_{t}+\beta_{\nu,9}\overline{\operatorname{RES}}_{t}$ (22)
$+\beta_{\nu,10}\operatorname{EUA}_{t}+\beta_{\nu,11}\operatorname{Gas}_{t}+\beta_{\nu,12}\operatorname{Coal}_{t}+\beta_{\nu,13}\operatorname{Oil}_{t}.$
@@
Scoring Rules
@@
We employ four well-established multivariate probabilistic scoring scores: The Energy Score (ES), the Dawid-Sebastiani Score (DSS), the Variogram Score (VS) and the Log-Score (LS). Additionally, we employ the root mean square error RMSE (RMSE), mean absolute error (MAE) and the continuous ranked probability score (CRPS). We test for statistically significant score differences using the well-established Diebold-Mariano test. The following paragraphs introduce the scores and are largely based on Gneiting et al. (2007); Gneiting and Raftery (2007); Nowotarski and Weron (2018); Marcotte et al. (2023); Ziel and Berk (2019) as well as the references mentioned for the individual scores. Denote the true price vector as $\bm{\mathrm{Y}}=(\bm{\mathrm{y}}_{0},...,\bm{\mathrm{y}}_{H})$ of shape $T\times H$ and an ensemble forecast $\bm{\mathrm{F}}$ of shape $T\times H\times M$ of $M=2500$ samples. The RMSE is defined as
@@
$\operatorname{CRPS}_{t}=\frac{1}{M}\sum_{m=0}^{M}|\bm{\mathrm{F}}_{t,h,m}-y_{t.h}|+\frac{1}{M}\sum_{m=0}^{M}\bm{\mathrm{F}}_{t,h,m}+\frac{1}{M(M-1)}\sum_{m=0}^{M}m\bm{\mathrm{F}}_{t,h,m}$ (27)
@@
The energy score is a strictly proper scoring rule, however, Alexander et al. (2024); Pinson and Tastu (2013); Marcotte et al. (2023) argue that the ES is rather insensitive to misspecified dependence structures. We aggregate the ES by taking the average: $\operatorname{ES}=\frac{1}{T}\sum_{t=0}^{T}\operatorname{ES}_{t}$. The Log-Score (LS) is defined as
$\operatorname{LS}_{t}=-\log\left(\mathcal{L}(\bm{\mathrm{y}}_{t}\mid\hat{\bm{\mathrm{\theta}}}^{\mathcal{D}}_{t})\right)$ (29)
@@
$\operatorname{DSS}_{t}=\log\left(\det(\widehat{\bm{\mathrm{\Sigma}}}_{F})\right)+(\bm{\mathrm{y_{t}}}-\bm{\mathrm{\hat{\mu}_{t}}})\widehat{\bm{\mathrm{\Sigma}}}_{F}^{-1}(\bm{\mathrm{y_{t}}}-\bm{\mathrm{\hat{\mu}_{t}}})$ (30)
@@
and is a proper scoring rule. We aggregate the VS by taking the average and normalize the score by dividing by $H^{2}$ and taking the square root, i.e. $\operatorname{VS}=\frac{1}{T}\sum_{t=0}^{T}\sqrt{\frac{1}{H^{2}}\operatorname{VS}_{t}}$ to make the scales of the score comparable. The scoring rules used are implemented in the Python package scoringrules (Zanetta and Allen, 2024).
Diebold-Mariano-Test
Conclusions on the performance of forecasting models cannot be drawn from looking at aggregate scores alone, but need to be drawn by evaluating whether the differential between the loss series of two models is statistically significantly from zero (Diebold and Mariano, 2002; Diebold, 2015). For the DM-test, we evaluate the differential of two score series $\Delta\bm{\mathrm{s}}^{A,B}=\bm{\mathrm{s}}^{A}-\bm{\mathrm{s}}^{B}$, where $\bm{\mathrm{s}}^{\mathcal{A}}=(s_{0}^{A},...,s_{T}^{A})$ are the scores for each scoring rule at $t$ for model $A$ respectively $B$. We provide two one-sided and hence complimentary tests.
@@ 4 Results
This section describes the results from the forecasting study. Exemplary simulations and predicted covariance matrices are shown in Figures 7 and 8. Some time-varying behavior of the covariance matrix over the week, especially in the morning hours is visible. We present aggregate scoring rules in Table 4 and significance testing using the DM-test in Figure 9 and discussed in the following paragraph. Computation times are discussed in the last paragraph and given in Table 5.

### v2

@@ 3.2 Model Definition and Benchmarks
We propose modeling the multivariate distribution of the day-ahead electricity prices in increasing complexity. If possible, models are updated online, i.e. using only the new data for each day. We differentiate between an adaptive estimation, which is updating a single, unconditional distributional parameter and the full conditional estimation linking the distribution parameter to explanatory variables.
@@
Figure 7: Model Taxonomy. We differentiate between modelling the marginal distribution and the dependence structure.
@@
Figure 7 shows the increasing model complexity. We start with the established LARX models and naively estimate the unconditional residual distribution (denoted as $\operatorname{LARX}+\mathcal{N}(0,\sigma)$ and $\operatorname{LARX}+\mathcal{N}(0,\Sigma)$). Additionally, we employ two additional univariate benchmarks. First, we use a conformal prediction (CP) approach, which is based on the LARX model and uses a combination of the split-conformal prediction (SCP) and adaptive conformal prediction (ACI, see Gibbs and Candes, 2021; Gibbs and Candès, 2024). We use absolute residual scores and a calibration set of 200 data points. Note that conformal prediction is generally a univariate post-processing method, since it is centered on issuing prediction intervals and therefore not suitable for the generation of ensemble forecasts. Multivariate conformal prediction approaches exists, but are in their infancy and focussed on a notion of multivariate quantiles. [^3] As a simple workaround, we have tried to combine CP with a copula-based approach using a PIT transformation. However, the conformal predictive density approximated using 199 quantiles was not sufficient to generate useful ensembles in the inverse transformation, i.e. the step from the simulated copula on $\mathcal{U}(0,1)$ to the original domain failed. Second, we use a GARCH(1,1) model with a Gaussian distributional assumption (denoted as LARX+GARCH(1,1)). We estimated the GARCH model on the residuals of the LARX model using the arch package in Python (Sheppard et al., 2024). We increase complexity by moving to full distributional model for the marginals and adding an adaptive estimation of the dependence structure using the Gaussian copula (denoted as oDistReg+Copula). Lastly, we estimate the full multivariate distribution in a conditional way using the proposed multivariate online distributional regression approach (denoted as oMvDistReg($\mathcal{F}$, parameterization, method)). We describe the full model in the following and note that we additionally describe the hyper parameters in A.5. For all three complexity levels, we include a reference model that assumes independence to showcase the value-add of modelling the dependence structure. We model the mean/location for all regression models, also for the LARX models , by
$g_{\mu}(\mu_{t,h})$ $=\beta_{\mu,0,h}+\sum_{l=1}^{L=7}\beta_{\mu,l,h}y_{t-l,h}+\sum_{i\in\{0,...,23\}\setminus h}\beta_{\mu,8+i,h}y_{t-1,i}+\sum_{w=1}^{W=6}\beta_{\mu,30+w,h}\operatorname{WD}_{t,h}$ (15)
$+\beta_{\mu,37,h}\operatorname{min}(\bm{\mathrm{y}}_{t-1})+\beta_{\mu,38,h}\operatorname{max}(\bm{\mathrm{y}}_{t-1})+\beta_{\mu,39,h}\operatorname{Q}_{10}(\bm{\mathrm{y}}_{t-1})+\beta_{\mu,40,h}\operatorname{Q}_{90}(\bm{\mathrm{y}}_{t-1})$
$+\beta_{\mu,41,h}\operatorname{ResLoad}_{t,h}+\beta_{\mu,42,h}\operatorname{EUA}_{t}+\beta_{\mu,43,h}\operatorname{Gas}_{t}+\beta_{\mu,44,h}\operatorname{Coal}_{t}+\beta_{\mu,45,h}\operatorname{Oil}_{t}.$
@@
$g_{\theta}(\theta_{t,h,h})$ $=\beta_{\theta,0,h}+\beta_{\theta,1,h}\operatorname{SignedSquare}\left(\bm{\mathrm{\Sigma}}_{h,h}^{[t-1:t-7]}\right)^{-1}+\beta_{\theta,2,h}\operatorname{ResLoad}_{t,h}$ (16)
$+\beta_{\theta,4,h}\operatorname{EUA}_{t}+\beta_{\theta,5,h}\operatorname{Gas}_{t}+\beta_{\theta,6,h}\operatorname{Coal}_{t}+\beta_{\theta,7,h}\operatorname{Oil}_{t},$
where $\operatorname{SignedSquare}(a)=\operatorname{sign}(a)\sqrt{|a|}$ is the signed square root and $\bm{\mathrm{\Sigma}}^{[t-1:t-7]}$ is the rolling empirical covariance matrix of $\bm{\mathrm{y}}_{t}$ for the last 7 days. Note that we use the inverse of the signed square, since we are working on the preicision matrix. For the univariate models, we replace this accordingly with the empirical rolling standard deviation. For the LRA-based parameterization, we choose $r=2$ and model the elements of $\bm{\mathrm{V}}$ as
$g_{v}(v_{t,h,0})$ $=\beta_{v,0,h}+\beta_{v,1,h}\operatorname{SignedSquare}\left(\bm{\mathrm{\Sigma}}_{h,h}^{[t-1:t-7]}\right)+\beta_{v,2,h}\overline{\operatorname{ResLoad}}_{t}$ (17)
$+\beta_{v,3,h}\operatorname{EUA}_{t}+\beta_{v,4,h}\operatorname{Gas}_{t}+\beta_{v,5,h}\operatorname{Coal}_{t}+\beta_{v,6,h}\operatorname{Oil}_{t},$
@@
$g_{\nu}(\nu_{t})$ $=\beta_{\nu,0}+\beta_{\nu,1}\operatorname{mean}(\bm{\mathrm{y}}_{t-1})+\sum_{w}^{W=6}\beta_{\nu,1+w,h}\operatorname{WD}_{t,h}+\beta_{\nu,8}\overline{\operatorname{ResLoad}}_{t}$ (19)
$+\beta_{\nu,9}\operatorname{EUA}_{t}+\beta_{\nu,10}\operatorname{Gas}_{t}+\beta_{\nu,11}\operatorname{Coal}_{t}+\beta_{\nu,12}\operatorname{Oil}_{t}.$
@@
3.3 Forecast Evaluation and Scoring Rules
Forecast evaluation should follow the well-known principle of sharpness subject to calibration Gneiting et al. (2007). We check the calibration of the forecasts by calculating joint prediction bands (JPB) from the simulations. JPB aim to cover the true price vector with certain probability $1-\alpha$ and can be thought of a multivariate generalization of marginal prediction intervals (Staszewska-Bystrova, 2011; Lütkepohl et al., 2015). JPBs have been used in energy market forecasting by Serafin et al. (2022); Chen et al. (2025). As it is standard in forecasting probabilistic forecasting, we evaluate both marginal and multivariate quality of the forecasts. Therefore, we employ the root mean square error RMSE (RMSE), mean absolute error (MAE) and the continuous ranked probability score (CRPS), which focus on mean/point prediction and the marginal distribution. We employ four well-established multivariate probabilistic scoring scores: The Energy Score (ES), the Dawid-Sebastiani Score (DSS), the Variogram Score (VS) and the Log-Score (LS). Let us briefly review some recent results with respect to the four multivariate scores:
  • 1.
    The ES, DSS and LS are all able to reliably detect mis-specifications in the mean structure of the multivariate distribution (Marcotte et al., 2023), while the VS (by design) is not sensitive. However, Pinson and Tastu (2013); Alexander et al. (2024) discuss the discrimination ability of the ES. While widely used for multivariate forecast evaluation, the ES has been shown to have low discrimination ability with respect to misspecified covariance structures, especially in double-digit and larger dimensions and Alexander et al. (2024) recommend to use the VS in addition to the ES.
  • 2.
    In a similar vein, Marcotte et al. (2023) show that the ES has low reliability, i.e. statistical power to discriminate between a correctly and incorrectly specified dependence model for multivariate forecasts compared to the DSS and the LS, with the VS being somewhat in the middle. Their analysis also reflects the role of the number of test samples (which, with more than 1800 out-of-sample days, is not a concern in our study) and the number of sample paths $M$. Interestingly, they find that the ES and the VS are complimentary, i.e. the VS is more reliable in cases where the ES is not and vice versa, and they find the reliability of the VS to increase non-monotonically with respect to the dimension $D$ and the number of samples $M$.
  • 3.
    In contrast, Ziel and Berk (2019) find that the ES, in combination with the Diebold-Marino test, is able to detect the correct model specification in a simulation study. Nevertheless, they still recommend the use of multiple scores.
@@ 3.3 Forecast Evaluation and Scoring Rules
Additionally, we test for statistically significant score differences using the well-established Diebold-Mariano test. The following paragraphs introduce JPBs and the scoring rules and are largely based on Gneiting et al. (2007); Gneiting and Raftery (2007); Nowotarski and Weron (2018); Marcotte et al. (2023); Ziel and Berk (2019) as well as the references mentioned for the individual scores. Denote the true price vector as $\bm{\mathrm{Y}}=(\bm{\mathrm{y}}_{0},...,\bm{\mathrm{y}}_{H})$ of shape $T\times H$ and the ensemble forecast as $\bm{\mathrm{F}}$ of shape $T\times H\times M$ of $M=2500$ samples.
A joint prediction band for the $1-\alpha$ coverage is defined by the lower bound $\bm{\mathrm{l}}_{t}=(l_{t,0},...,l_{t,23})$ and upper bound $\bm{\mathrm{u}}_{t}=(u_{t,0},...,u_{t,23})$, such that
$\mathbb{P}_{t}\left(\bm{\mathrm{l}}_{t}\leq\bm{\mathrm{y}}_{t}\leq\bm{\mathrm{u}}_{t}\;\forall\;h\in\{0,...,23\}\right)=1-\alpha,$
that is, the true price trajectory $\bm{\mathrm{y}}_{t}$ is fully covered by the JPB with probability $1-\alpha$. This is in difference to marginal prediction intervals based on predicted quantiles, which consider element-wise coverage. The difference between marginal, quantile-based prediction bands and joint prediction bands is shown in Figure 8. There are multiple algorithms to construct such bands and following Serafin et al. (2022); Chen et al. (2025), we use the neighouring paths method described by Staszewska-Bystrova (2011); Lütkepohl et al. (2015). The methods is based on iteratively removing paths from the ensemble forecast $\bm{\mathrm{F}}$, such that the envelope of the remaining paths covers the true price trajectory $\bm{\mathrm{y}}_{t}$ with probability $1-\alpha$. It should be noted that JPB are in general not unique, i.e. there are multiple sets of lower and upper bounds that cover the true trajectory with the desired probability. We compare the mis-coverage of the prediction interval, which is defined as
$\operatorname{MC}_{1-\alpha}=\frac{1}{T}\sum^{T}_{t=0}\bm{1}\left(\bm{\mathrm{l}}_{t}\leq\bm{\mathrm{y}}_{t}\leq\bm{\mathrm{u}}_{t}\;\forall\;h\in\{0,...,23\}\right)-(1-\alpha),$ (22)
and gives a measure of the multivariate calibration of the joint prediction. Additionally, we evaluate the mean width of the prediction bands, which is defined as
$\operatorname{JPBW}_{1-\alpha}=\frac{1}{TH}\sum^{T}_{t=0}\sum^{H}_{h=0}\left(\bm{\mathrm{u}}_{t,h}-\bm{\mathrm{l}}_{t,h}\right),$ (23)
which gives a measure of the efficiency of the prediction bands and can be interpreted as the area of the prediction band divided by the dimensions.
Figure 8: Illustrative 50%-prediction bands based on marginal quantiles and the iterative method for JPBs presented in Lütkepohl et al. (2015) and Staszewska-Bystrova (2011). In the upper panel, the simulations disregard any dependece between the 24 delivery hours, otherwise the marginal distributions are identical. Hence, the quantile prediction bands are similar (up to simulation noise), while for the joint prediction bands, modelling the dependence decreases the width of the prediction bands. Own illustration.
The RMSE is defined as
@@
$\operatorname{CRPS}_{t}=\frac{1}{M}\sum_{m=0}^{M}|\bm{\mathrm{F}}_{t,h,m}-y_{t.h}|+\frac{1}{M}\sum_{m=0}^{M}\bm{\mathrm{F}}_{t,h,m}+\frac{1}{M(M-1)}\sum_{m=0}^{M}m\bm{\mathrm{F}}_{t,h,m}.$ (26)
@@
The energy score is a strictly proper scoring rule. We aggregate the ES by taking the average: $\operatorname{ES}=\frac{1}{T}\sum_{t=0}^{T}\operatorname{ES}_{t}$. The Log-Score (LS) is defined as
$\operatorname{LS}_{t}=-\log\left(\mathcal{L}(\bm{\mathrm{y}}_{t}\mid\hat{\bm{\mathrm{\theta}}}^{\mathcal{D}}_{t})\right),$ (28)
@@
$\operatorname{DSS}_{t}=\log\left(\det(\widehat{\bm{\mathrm{\Sigma}}}_{F})\right)+(\bm{\mathrm{y_{t}}}-\bm{\mathrm{\hat{\mu}_{t}}})\widehat{\bm{\mathrm{\Sigma}}}_{F}^{-1}(\bm{\mathrm{y_{t}}}-\bm{\mathrm{\hat{\mu}_{t}}}),$ (29)
@@
and is a proper scoring rule. We aggregate the VS by taking the average and normalize the score by dividing by $H^{2}$, i.e. $\operatorname{VS}=\frac{1}{TH^{2}}\sum_{t=0}^{T}{\operatorname{VS}_{t}}$ to make the scales of the score comparable. The scoring rules used are implemented in the Python package scoringrules (Zanetta and Allen, 2024).
Conclusions on the performance of forecasting models cannot be derived by looking at aggregate scores alone, but need to be drawn by evaluating whether the differential between the loss series of two models is statistically significantly from zero (Diebold and Mariano, 2002; Diebold, 2015). For the DM-test, we evaluate the differential of two score series $\Delta\bm{\mathrm{s}}^{\mathcal{A},\mathcal{B}}=\bm{\mathrm{s}}^{\mathcal{A}}-\bm{\mathrm{s}}^{\mathcal{B}}$, where $\bm{\mathrm{s}}^{\mathcal{A}}=(s_{0}^{\mathcal{A}},...,s_{T}^{\mathcal{A}})$ are the scores for each scoring rule at $t$ for model $\mathcal{A}$ respectively $\mathcal{B}$. We provide two one-sided and hence complimentary tests. To ensure validity of the DM-test, we check stationarity of the differential series $\Delta\bm{\mathrm{s}}^{\mathcal{A},\mathcal{B}}$ by the augmented Dickey-Fuller test (ADF, Dickey and Fuller, 1979; Cheung and Lai, 1995).
@@ 4 Results
We present the results of our forecasting study in three parts. First, we illustrate the differences between the models by showing example simulations and analyzing the time-varying dependence structure. Second, we present the forecasting accuracy of all models using the scoring rules presented above. Third, we analyze the role of overfitting and regularization in our proposed path-based estimation approach. Lastly, we discuss the computational costs of the different models compare the online and batch estimation. Figure 9 shows illustrative simulations drawn from two models. Figure 10 shows exemplary predicted covariance matrices.
Figure 9: Illustrative simulations drawn from two models.